# Qini Curve Estimation Under Clustered Network Interference

### v1

@@ Abstract
Qini curves are a widely used tool for assessing treatment policies under allocation constraints as they visualize the incremental gain of a new treatment policy versus the cost of its implementation. Standard Qini curve estimation assumes no interference between units: that is, that treating one unit does not influence the outcome of any other unit. In many real-life applications such as public policy or marketing, however, the presence of interference is common. Ignoring interference in these scenarios can lead to systematically biased Qini curves that over- or under-estimate a treatment policy’s cost-effectiveness. In this paper, we address the problem of Qini curve estimation under clustered network interference, where interfering units form independent clusters. We propose a formal description of the problem setting with an experimental study design under which we can account for clustered network interference. Within this framework, we introduce three different estimation strategies suited for different conditions. Moreover, we introduce a marketplace simulator that emulates clustered network interference in a typical e-commerce setting. From both theoretical and empirical insights, we provide recommendations in choosing the best estimation strategy by identifying an inherent bias-variance trade-off among the estimation strategies.
**footnotetext: Equal contribution.
1 Introduction
Understanding treatment effect heterogeneity –the variation in individual responses to the same treatment within a population– is central in shaping individualized treatment policies across various domains, including personalized medicine (Kravitz et al., 2004), uplift modeling in marketing and e-commerce (Goldenberg et al., 2020), and targeted subgroup interventions in public policy (Brand and Davis, 2011). In these scenarios, the same questions recurs: Who should we treat? Sometimes, it is sufficient to identify individuals who respond positively to a treatment. However, when treatments involve monetary or practical costs, the challenge is to devise a cost-effective policy that targets those who benefit the most from the treatment while staying within a given budget for treatment allocation.
@@ 1 Introduction
Interference arises in a variety of contexts, from peer effects in social networks (Manski, 2013; Ogburn et al., 2020) to cannibalization effects on marketplace platforms  (Holtz et al., 2024). One of the most common settings is so-called clustered network interference where interference only happens within, rather than between, clusters. While there exists an extensive body of literature on estimating treatment effects under clustered network interference, e.g. Sobel (2006); Hudgens and Halloran (2008), little attention has been given to the problem of estimating Qini curves in this setting. As we demonstrate in Figure 1, traditional methods for estimating Qini curves become significantly biased when interference is present. Since biased Qini curves lead to incorrect assessments of the cost-effectiveness of treatment policies, this is an important yet unaddressed problem. Consequently, the central question we aim to answer in this paper is: How can we accurately estimate Qini curves under clustered network interference?
@@
To address our research question, we first present a formal description of the problem and describe the experimental study design and necessary identification conditions for estimating Qini curves under clustered network interference. Next, we propose three different estimation strategies based on different modeling assumptions, highlighting their respective benefits and drawbacks from a bias-variance perspective. We then empirically compare our proposed methods with a traditional approach for Qini curve estimation using a simulated dataset designed to mimic a marketplace with interference in the form of cannibalization among different vendors. Finally, based on our findings, we provide practical recommendations on how to estimate Qini curves in settings with clustered network interference.
Figure 1: An illustrative experiment with Qini curve estimation under clustered network interference. The black dashed line represents the true underlying Qini curve, while the solid lines depict two estimation approaches: one based on a traditional method that assumes no interference, and the other representing a proposed strategy in this paper that adjusts for interference using inverse probability weighting. More details on the simulation used to generate this figure can be found in Appendix C.
@@ 2 Related works
The task of estimating treatment effects becomes considerably more complex in the presence of interference. Hence, despite early influential works in causal inference such as Rubin (1974), only recently has a large body of literature emerged to tackle scenarios with interference. One of the most commonly studied settings is clustered network interference, where treatment units form independent clusters (Sobel, 2006; Hudgens and Halloran, 2008; Tchetgen and VanderWeele, 2012), a condition also known as partial interference. Interference also naturally arises in network data, where some units are related to other units by being neighbors in e.g. a social network (Ugander et al., 2013; Eckles et al., 2017). In some cases, an experimental study design can be constructed in a way to detect and reduce bias from interference, for instance, through a two-stage randomization design (Hudgens and Halloran, 2008) or by stratified randomization across different blocks (Bajari et al., 2021). Previous works have covered specific tasks under interference such as heterogeneous treatment effect estimation (Zhao et al., 2024) or policy evaluation/learning (Zhang and Imai, 2023). To our knowledge, there is no prior work on the problem of estimating Qini curves in the presence of interference.
Evaluating treatment prioritization rules using Qini curves in settings without interference has gained more attention in recent years (Radcliffe, 2007; Rößler and Schoder, 2022; Bokelmann and Lessmann, 2024). The development of estimations procedures with better statistical inference guarantees has enabled the use of Qini curves in this context (Yadlowsky et al., 2024). While none of these works consider interference, Sverdrup et al. (2024) considers the related problem of estimating Qini curves for combinatorial multi-armed treatments. There is an inherent connection between combinatorial treatment problems and clustered network interference, as the treatment assignment of units within a single cluster can be seen as a combinatorial treatment decision. This connection also underscores the challenge of estimating Qini curves under clustered network interference: as cluster size grows the combinatorial space of possible treatments expands exponentially, leading to a corresponding increase in interactions among units within the cluster. To address this challenge, we propose estimation strategies designed to more accurately estimate Qini curves, even as the cluster size grows.
@@ 3 Data structure & assumptions
We assume access to observations from a distribution $P$. We have clusters $i=1,\dots,N$ and each cluster contains the units $j=1,\dots,M_{i}$. A unit can be referred to by the tuple $(i,j)$. For each cluster $i$, we observe pre-treatment covariates $X_{i}$ in $\mathcal{X}\subseteq\mathbb{R}^{d_{x}}$. For each unit, we observe pre-treatment covariates $Z_{ij}$ in $\mathcal{Z}\subseteq\mathbb{R}^{d_{z}}$, a binary treatment $W_{ij}\in\{0,1\}$, and an outcome of interest $Y_{ij}$ in $\mathcal{Y}\subseteq\mathbb{R}$. The outcome may be binary or continuous. In addition, we also observe a non-negative cost of treatment $C_{ij}$ in $\mathcal{C}\subseteq[0,\infty)$. The cost $C_{ij}$ depends on both the treatment and outcome, and specifically we assume there to be no cost $C_{ij}=0$ when no treatment is given $W_{ij}=0$. We consider cluster-level outcomes and costs which we define as ${Y}_{i}=\sum_{j=1}^{M_{i}}Y_{ij}$ and ${C}_{i}=\sum_{j=1}^{M_{i}}C_{ij}$. We also define the cluster-level treatment which is a binary vector $\mathbf{W}_{i}=[W_{i1},W_{i2},\dots,W_{iM_{i}}]\in\{0,1\}^{M_{i}}$. At last, random variables are denoted by capital letters, while their instantiated values use lowercase. Probability densities are represented as $f(\cdot)$.
@@
In the setting of clustered network interference, we assume observations can be divided in independent clusters. Treating one unit belonging to cluster $i$ may influence the outcomes of other units from that same cluster. However, treating units from a different cluster $i^{\prime}$ will not influence the outcomes of the units in cluster $i$. To define causal effects in this setting, we posit potential (counterfactual) outcomes $Y_{ij}(\mathbf{w})$ corresponding to the outcomes we would observe for an unit$(i,j)$ if the treatment vector $\mathbf{W}_{i}$ would be set to $\mathbf{w}$ (Tchetgen and VanderWeele, 2012). Analogously, we define the counterfactual cost $C_{ij}(\mathbf{w})$ if $\mathbf{W}_{i}$ would be set to $\mathbf{w}$.
@@
Throughout this paper, we consider an experimental study design where the unit-level treatments $W_{ij}$ are independently and randomly assigned. The treatment probability is determined by $e_{w}(x)=\Pr(W=w\mid X=x)$ which is known and the same for all units within a given cluster. Following standard convention, we refer to this probability as the propensity score (Rosenbaum and Rubin, 1983). We assume the following conditions are fulfilled by our experimental study design.
Assumption 3.1.
@@
Consistency is met when the intervention is unambiguously defined, meaning that no undisclosed variants of the treatment exist. Conditional exchangeability corresponds assuming no unmeasured confounding; specifically, the characteristics captured by cluster-level covariates are sufficient to control for any confounding between treatment assignment and outcome/cost. Positivity necessitates that all clusters have a non-zero probability of receiving any of combination of available treatments among its units. In the context of our experimental study design, we emphasize that conditional exchangeability and positivity can be guaranteed by (conditional) randomization.
@@ 4 Assessing treatment policies using Qini curves under clustered network interference
We are interested in assessing treatment policies based on some treatment prioritization rule $S:\mathcal{X}\times\mathcal{Z}\rightarrow\mathbb{R}$ that attempts to rank all units across the clusters based on who responds best to the treatment. A larger $S(X_{i},Z_{ij})$ should here be interpreted as that the unit $j$ in cluster $i$ is expected to have a larger treatment effect. Given a treatment prioritization rule $S$ and a fixed treatment threshold $R\in\mathbb{R}$, we will evaluate decisions by treatment policies defined as
$\pi_{S,R}(x,z)=\begin{cases}1,&S(x,z)\geq R\\ 0,&S(x,z)<R\end{cases}~{}.$
@@ 4.1 Definition of the Qini curve
A common approach to assess how well a treatment prioritization rule identifies those who best respond to a treatment is by using Qini curves (Radcliffe, 2007; Sverdrup et al., 2024). We denote the decisions made by the policy $\pi$ on an evaluation dataset as $\pi_{ij}=\pi(X_{i},Z_{ij})\in\{0,1\}$ for unit $(i,j)$ and $\boldsymbol{\pi}_{i}=[\pi(X_{i},Z_{i1}),\dots,\pi(X_{i},Z_{iM_{i}})]\in\{0,1\}^{M_{i}}$ for the collective treatment decision on cluster $i$. Then, we first define the policy value in terms of average cluster-level outcomes under decisions made by the policy $\pi$,
$V(\pi)=\mathbb{E}\left[\sum_{j=1}^{M_{i}}Y_{ij}\left(\boldsymbol{\pi}_{i}\right)\right]~{},$
@@
$C(\pi)=\mathbb{E}\left[\sum_{j=1}^{M_{i}}C_{ij}\left(\boldsymbol{\pi}_{i}\right)\right]~{}.$
The expectations are taken with respect to the distribution of all possible clusters.
Let $R_{B}$ denote the threshold such that $C(\pi_{S,R_{B}})=B$, then we can define the Qini curve for a treatment prioritization rule $S$ as follows:
$Q_{S}(B)=V(\pi_{S,R_{B}})-V(\pi_{0}),\;\;B\in[0,B_{\text{max}}]$ (1)
where $\pi_{0}\equiv 0$ is a reference policy that treats none and $B_{\text{max}}>0$ is the maximal allowed cost under consideration.
The above definition is more general than the one by Radcliffe (2007), who assumes uniform cost across all units. In their approach, $Q_{S}(B)$ on the y-axis against the fraction of treated units on the x-axis. The definition presented here can be applied to this case by plotting $B/B_{\text{max}}$ on the x-axis, where $B_{\text{max}}$ represents the total cost of treating all units. We refer to this as the uniform cost case.
To understand how interference introduces challenges in the estimation of Qini curves, consider a scenario where two units from the same cluster fall on opposite sides of the threshold $R$ –one above (indicating it should be treated) and one below (indicating it should not). Normally, these units would be considered independent, but in the presence of interference, spillover effects may occur between them. If these effects are not accounted for, we might over- or underestimate the policy value and cost which also affects the Qini curve estimation.
For the remainder of this paper, we will demonstrate how to address this issue and provide a methodology for estimating Qini curves that appropriately accounts for interference within a clustered network setting. However, before we discuss estimation, we also establish a necessary identifiability result that allows for estimation.
@@ 4.2 Identifiability of policy value and policy cost
To estimate $Q_{S}(B)$ in (1), $V(\pi)$ and $C(\pi)$ must be identifiable from the observed data. We note that under assumption 3.1, these can be identified from the data collected in our randomized design. More specifically, recall that $Y_{i}=\sum_{j=1}^{M_{i}}Y_{ij}$ and $C_{i}=\sum_{j=1}^{M_{i}}C_{ij}$, then we define
$\phi(\pi)$ $=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}[Y_{i}\mid\mathbf{W}=\boldsymbol{\pi}_{i},X_{i}]~{},$
$\psi(\pi)$ $=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}[C_{i}\mid\mathbf{W}=\boldsymbol{\pi}_{i},X_{i}]~{},$
for which we can show that the policy value and policy cost are identifiable from the observed data (see Appendix  A.1 for the proof).
Theorem 4.1.
Under assumption 3.1, we have that $V(\pi)=\mathbb{E}[\phi(\pi)]$ and $C(\pi)=\mathbb{E}[\psi(\pi)]$.
@@
0:  Dataset $D=\{X_{i},\{Z_{ij},W_{ij},Y_{ij},C_{ij}\}_{j=1}^{M_{i}}\}_{i=1}^{N}$; treatment prioritization rule $S$; number of percentiles $K$; max budget $B_{\text{max}}$; estimators $\widehat{\phi}$ and $\widehat{\psi}$; Boolean flag indicating if cost is uniform
1:  Set $\widehat{V}_{0}=\widehat{\phi}(\pi_{0}\equiv 0)$ and $(\widehat{B}_{0},\widehat{Q}_{0})=(0,0)$
2:  Let $S_{\text{sorted}}(i)$ be the score for the $i$th unit when sorted by $S$ in descending order
3:  for $k\in[1,\dots,K]$ do
4:     Set $R_{B}=S_{\text{sorted}}(i_{k})$ where $i_{k}=\text{round}(\frac{k}{K}\cdot|D|$) and $|D|$ is the total number of units
5:     if cost is uniform then
6:        Set $\widehat{B}_{k}=\frac{k}{K}\cdot B_{\text{max}}$
7:     else
8:        Set $\widehat{B}_{k}=\widehat{\psi}(\pi_{S,R_{B}})$
9:     end if
10:     Set $\widehat{Q}_{k}=\widehat{\phi}(\pi_{S,R_{B}})-\widehat{V}_{0}$
11:  end for
12:  return $\{\widehat{B}_{k},\widehat{Q}_{k}\}_{k=0}^{K}$
With the established identification results, we can now propose a general procedure for Qini curve estimation in our experimental study design. In practice, to estimate the Qini curve for a treatment prioritization rule $S$, one typically estimates the policy value and cost over a range of pre-specified thresholds $R$. For the uniform cost case, it is sufficient to only estimate the policy value. So far we assume ranking according to $S$ leads to no ties, but if there are ties one could add tiebreakers depending on their application, e.g., a small amount of random noise.
The full procedure is shown in Algorithm 1. To implement this algorithm, we need strategies for estimating $\phi(\pi)$ and $\psi(\pi)$, which will be the focus of the coming section.

### v2

@@ Abstract
Qini curves are a widely used tool for assessing treatment policies under allocation constraints as they visualize the incremental gain of a new treatment policy versus the cost of its implementation. Standard Qini curve estimation assumes no interference between units: that is, that treating one unit does not influence the outcome of any other unit. In many real-life applications such as public policy or marketing, however, the presence of interference is common. Ignoring interference in these scenarios can lead to systematically biased Qini curves that over- or under-estimate a treatment policy’s cost-effectiveness. In this paper, we address the problem of Qini curve estimation under clustered network interference, where interfering units form independent clusters. We propose a formal description of the problem setting with an experimental study design under which we can account for clustered network interference. Within this framework, we describe three estimation strategies, each suited to different conditions, and provide guidance for selecting the most appropriate approach by highlighting the inherent bias-variance trade-offs. To complement our theoretical analysis, we introduce a marketplace simulator that replicates clustered network interference in a typical e-commerce environment, allowing us to evaluate and compare the proposed strategies in practice.
Understanding treatment effect heterogeneity – the variation in individual responses to the same treatment within a population – is central in shaping individualized treatment policies across various domains, including personalized medicine (Kravitz et al., 2004), uplift modeling in marketing and e-commerce (Goldenberg et al., 2020), and targeted subgroup interventions in public policy (Brand & Davis, 2011). In these scenarios, the same questions recurs: Who should we treat? Sometimes, it is sufficient to identify individuals who respond positively to a treatment. However, when treatments involve monetary or practical costs, the challenge is to devise a cost-effective policy that targets those who benefit the most from the treatment while staying within a given budget for treatment allocation.
@@
Interference arises in a variety of contexts, from peer effects in social networks (Manski, 2013; Ogburn et al., 2020) to cannibalization effects on marketplace platforms  (Holtz et al., 2024). One of the most common settings is so-called clustered network interference where interfering units form independent clusters; here units interfere within, rather than between, clusters. While there exists an extensive body of literature on estimating treatment effects under clustered network interference, e.g. Sobel (2006); Hudgens & Halloran (2008), little attention has been given to the problem of estimating Qini curves in this setting. As we demonstrate in Figure 1, traditional methods for estimating Qini curves become significantly biased when interference is present. Since biased Qini curves lead to incorrect assessments of the cost-effectiveness of treatment policies, this is an important yet unaddressed problem. For example, Imai & Li (2023) emphasize the need for methods that evaluate individualized treatment strategies under interference. Motivated by this gap, the central question we aim to answer in this paper is: How can we account for interference when estimating Qini curves, and how does this adjustment affect decision-making using Qini curves?
@@
To address our research question, we first formulate the experimental study design and necessary identification conditions for estimating Qini curves under clustered network interference. Next, we describe three separate estimation strategies based on different assumptions of the underlying interference. Our theoretical analysis demonstrates that stronger assumptions on the interference can yield more efficient estimators, though at the potential cost of increased bias when those assumptions are violated. We explore these trade-offs by empirically comparing all methods using a simulated data-generating process designed to mimic a marketplace with different types of interference in the form of cannibalization among different vendors. Finally, based on our findings, we offer practical recommendations for estimating Qini curves in settings with clustered network interference.
Figure 1: An illustrative simulation study with Qini curve estimation under clustered network interference. The black dashed line represents the true underlying Qini curve, while the solid lines depict two estimation approaches: one based on a traditional method that assumes no interference, and the other representing a proposed strategy in this paper that adjusts for interference using inverse probability weighting. More details on the simulation used to generate this figure can be found in Appendix D.
@@ 1 Related works
The task of estimating treatment effects becomes considerably more complex in the presence of interference. Hence, despite early influential works in causal inference such as Rubin (1974), only recently has a substantial body of literature emerged to address interference. One of the most commonly studied settings is clustered network interference, where treatment units form independent clusters (Sobel, 2006; Hudgens & Halloran, 2008; Tchetgen & VanderWeele, 2012), a condition also known as partial interference. Interference also naturally arises in network data, where some units are related to other units by being neighbors in e.g. a social network (Ugander et al., 2013; Eckles et al., 2017). In some cases, an experimental study design can be constructed in a way to detect and reduce bias from interference, for instance, through a two-stage randomization design (Hudgens & Halloran, 2008) or by stratified randomization across different blocks (Bajari et al., 2021). Prior work has tackled specific tasks under interference, such as heterogeneous treatment effect estimation (Zhao et al., 2024) and policy evaluation/learning (Zhang & Imai, 2025). To our knowledge, however, no prior work has considered the problem of estimating Qini curves in the presence of interference.
Evaluating treatment prioritization rules using Qini curves in settings without interference has gained more attention in recent years (Radcliffe, 2007; Rößler & Schoder, 2022; Bokelmann & Lessmann, 2024). The development of estimations procedures with better statistical inference guarantees has enabled the use of Qini curves in this context (Yadlowsky et al., 2024). While none of these works consider interference, Sverdrup et al. (2025) considers the related problem of estimating Qini curves for combinatorial multi-armed treatments.
Combinatorial treatment problems and clustered network interference are inherently connected, as assigning treatments within a cluster can be seen as a combinatorial treatment decision. While causal inference methods exist for such settings, see e.g. Dasgupta et al. (2014); Goplerud et al. (2025), they typically assume a fixed treatment dimension – an assumption too restrictive when cluster sizes also may vary. However, the connection between these high-dimensional treatment settings and our setting underscores the challenge of estimating Qini curves under clustered network interference: as cluster size grows the combinatorial space of possible treatments expands exponentially, leading to a corresponding increase in interactions among units within the cluster. To address this challenge, we propose estimation strategies designed to more accurately estimate Qini curves, even as the cluster size differs or grows.
@@ 2 Data structure & assumptions
We assume access to observations from a distribution $P$. We have clusters $i=1,\dots,N$ and each cluster contains the units $j=1,\dots,M_{i}$. A unit can be referred to by the tuple $(i,j)$. For each cluster $i$, we observe pre-treatment covariates $X_{i}$ in $\mathcal{X}\subseteq\mathbb{R}^{d_{x}}$. For each unit, we observe pre-treatment covariates $Z_{ij}$ in $\mathcal{Z}\subseteq\mathbb{R}^{d_{z}}$, a binary treatment $W_{ij}\in\{0,1\}$, and an outcome of interest $Y_{ij}$ in $\mathcal{Y}\subseteq\mathbb{R}$. The outcome may be binary or continuous. In addition, we also observe a non-negative cost of treatment $C_{ij}$ in $\mathcal{C}\subseteq[0,\infty)$. The cost $C_{ij}$ depends on both the treatment and outcome, and specifically we assume there to be no cost $C_{ij}=0$ when no treatment is given $W_{ij}=0$. We consider cluster-level outcomes and costs which we define as $\tilde{Y}_{i}=\sum_{j=1}^{M_{i}}Y_{ij}$ and $\tilde{C}_{i}=\sum_{j=1}^{M_{i}}C_{ij}$. We also define the cluster-level treatment which is a binary vector $\mathbf{W}_{i}=[W_{i1},W_{i2},\dots,W_{iM_{i}}]\in\{0,1\}^{M_{i}}$. At last, random variables are denoted by capital letters, while their instantiated values use lowercase. Probability densities are represented as $f(\cdot)$.
@@
In the setting of clustered network interference, we assume observations can be divided in independent clusters. Treating one unit belonging to cluster $i$ may influence the outcomes of other units from that same cluster. However, treating units from a different cluster $i^{\prime}$ will not influence the outcomes of the units in cluster $i$. To define causal effects in this setting, we posit potential (counterfactual) outcomes $Y_{ij}(\mathbf{w})$ corresponding to the outcomes we would observe for an unit $(i,j)$ if the treatment vector $\mathbf{W}_{i}$ would be set to $\mathbf{w}$ (Tchetgen & VanderWeele, 2012). Analogously, we define the counterfactual cost $C_{ij}(\mathbf{w})$ if $\mathbf{W}_{i}$ would be set to $\mathbf{w}$.
@@
Throughout this paper, we consider an experimental study design where the unit-level treatments $W_{ij}$ are independently and randomly assigned. The treatment probability is determined by $e_{w}(x)=\Pr(W=w\mid X=x)$ which is known and the same for all units within a cluster. Following standard convention, we refer to this probability as the propensity score (Rosenbaum & Rubin, 1983). We assume the following conditions are fulfilled by our experimental study design.
Assumption 1.
@@
Consistency is met when the intervention is unambiguously defined, meaning that no undisclosed variants of the treatment exist. Conditional exchangeability corresponds to assuming no unmeasured confounding; specifically, the characteristics captured by cluster-level covariates are sufficient to control for any confounding between treatment assignment and outcome/cost. Positivity necessitates that all clusters have a non-zero probability of receiving any of combination of available treatments among its units. In the context of our experimental study design, we emphasize that conditional exchangeability and positivity can be guaranteed by (conditional) randomization.
@@ 3 Assessing treatment policies using Qini curves under clustered network interference
We are interested in assessing treatment policies based on some treatment prioritization rule $S:\mathcal{X}\times\mathcal{Z}\rightarrow\mathbb{R}$ that attempts to rank all units across the clusters based on who responds best to the treatment. A larger $S(X_{i},Z_{ij})$ should here be interpreted as that the unit $(i,j)$ is expected to have a larger treatment effect. Given a treatment prioritization rule $S$ and a fixed treatment threshold $R\in\mathbb{R}$, we will evaluate decisions by treatment policies defined as
$\pi_{S,R}(x,z)=\begin{cases}1,&S(x,z)\geq R\\ 0,&S(x,z)<R\end{cases}~.$ (1)
@@
For clarity and consistency with prior work on Qini curve estimation, we present scoring rules that depend only on unit- and cluster-level covariates, ignoring interference effects. In principle, however, scoring rules could incorporate treatments and covariates of other units within the same cluster. Notably, the identifiability results at the end of this section remain valid under such generalized scoring rules, though policies based on them may be difficult to implement in practice.
@@ 3.1 Definition of the Qini curve
A common approach to assess how well a treatment prioritization rule identifies those who best respond to a treatment is by using Qini curves (Radcliffe, 2007; Sverdrup et al., 2025). We denote the decisions made by the policy $\pi$ on an evaluation dataset as $\pi_{ij}=\pi(X_{i},Z_{ij})\in\{0,1\}$ for unit $(i,j)$ and $\boldsymbol{\pi}_{i}=[\pi(X_{i},Z_{i1}),\dots,\pi(X_{i},Z_{iM_{i}})]\in\{0,1\}^{M_{i}}$ for the collective treatment decision on cluster $i$. Then, we first define the policy value in terms of average cluster-level outcomes under decisions made by the policy $\pi$ as
$V(\pi)=\mathbb{E}\left[\sum_{j=1}^{M_{i}}Y_{ij}\left(\boldsymbol{\pi}_{i}\right)\right]~,$ (2)
@@
$C(\pi)=\mathbb{E}\left[\sum_{j=1}^{M_{i}}C_{ij}\left(\boldsymbol{\pi}_{i}\right)\right]~.$ (3)
Let $R_{B}$ denote the treatment threshold such that $C(\pi_{S,R_{B}})=B$, then we can define the Qini curve for a treatment prioritization rule $S$ as follows:
$Q_{S}(B)=V(\pi_{S,R_{B}})-V(\pi_{0}),\;\;B\in[0,B_{\text{max}}]~,$ (4)
where $\pi_{0}\equiv 0$ is a reference policy that treats none and $B_{\text{max}}>0$ is the maximal allowed cost by the treatment strategy under consideration.
The above definition is more general than the one by Radcliffe (2007), who assumes uniform cost across all units. In their approach, one plots $Q_{S}(B)$ on the y-axis against the fraction of treated units on the x-axis. The definition presented here can be applied to this case by plotting $B/B_{\text{max}}$ on the x-axis, where $B_{\text{max}}$ represents the total cost of treating all units. We refer to this as the uniform cost case.
To understand how interference introduces challenges in the estimation of Qini curves, consider a scenario where two units from the same cluster fall on opposite sides of the treatment threshold $R$ – one above (indicating it should be treated) and one below (indicating it should not). Normally, these units would be considered independent, but in the presence of interference, spillover effects may occur between them. If these effects are not accounted for, we might over- or underestimate the policy value and cost which also affects the Qini curve estimation.
For the remainder of this paper, we will demonstrate how to address this issue and provide a methodology for estimating Qini curves that appropriately accounts for interference within a clustered network setting. However, before we propose multiple estimation strategies, we also establish a necessary identifiability result that enables the estimation.
@@ 3.2 Identifiability of policy value and policy cost
To estimate $Q_{S}(B)$ in (4), $V(\pi)$ and $C(\pi)$ must be identifiable from the observed data. Under a study design that satisfies Assumption 1, this identifiability is guaranteed. More specifically, recall that $\tilde{Y}_{i}=\sum_{j=1}^{M_{i}}Y_{ij}$ and $\tilde{C}_{i}=\sum_{j=1}^{M_{i}}C_{ij}$, then we define
$\phi(\pi)=\mathbb{E}\big[\mathbb{E}\big[\tilde{Y}\mid\mathbf{W}=\boldsymbol{\pi},X\big]\big]\quad\text{and}\quad\psi(\pi)=\mathbb{E}\big[\mathbb{E}\big[\tilde{C}\mid\mathbf{W}=\boldsymbol{\pi},X\big]\big]~.$ (5)
With these definitions, we obtain the following important identifiability result (see proof in Appendix A.1).
Lemma 1.
Under Assumption 1, we have $V(\pi)=\phi(\pi)$ and $C(\pi)=\psi(\pi)$.
With the established identification results, we can now outline the general procedure for Qini curve estimation in our experimental study design. In practice, to estimate the Qini curve for a treatment prioritization rule $S$, one typically estimates the policy value $\phi(\pi)$ and cost $\psi(\pi)$ over a range of pre-specified thresholds $R$. For the uniform cost case, it is sufficient to only estimate the policy value. We outline the full procedure in Algorithm 1. So far, we assumed ranking according to $S$ leads to no ties, but if there are ties one could add tiebreakers by, e.g., injecting a small amount of random noise. To implement this algorithm we would need estimators for $\phi(\pi)$ and $\psi(\pi)$; for this reason, next we describe three estimation strategies designed for different scenarios.
@@
1:Dataset $D=\{X_{i},\{Z_{ij},W_{ij},Y_{ij},C_{ij}\}_{j=1}^{M_{i}}\}_{i=1}^{N}$; treatment prioritization rule $S$; number of percentiles $K$; max budget $B_{\text{max}}$; estimators $\widehat{\phi}$ and $\widehat{\psi}$; Boolean flag indicating if cost is uniform
[…]
4:Let $S_{\text{sorted}}(i)$ be the score for the $i$th unit when sorted by $S$ in descending order
5:for $k=1$ to $K$ do
6:  $i_{k}\leftarrow\text{round}\!\left(\tfrac{k}{K}\cdot|D|\right)$
7:  $R_{B}\leftarrow S_{\text{sorted}}(i_{k})$
8:  if cost is uniform then
9:   $\widehat{B}_{k}\leftarrow\tfrac{k}{K}\cdot B_{\text{max}}$
10:  else
11:   $\widehat{B}_{k}\leftarrow\widehat{\psi}(\pi_{S,R_{B}})$
12:  end if
13:  $\widehat{Q}_{k}\leftarrow\widehat{\phi}(\pi_{S,R_{B}})-\widehat{V}_{0}$
14:end for
15:return $\{(\widehat{B}_{k},\widehat{Q}_{k})\}_{k=0}^{K}$